OpenStudy (anonymous):

The perimeter of a basketball court is 102 meters and the length is 6 meters longer than twice the width. what are the length and the width?

OpenStudy (midhun.madhu1987):

 Solve for x and find the length and width..

OpenStudy (anonymous):

thats my problem. im not sure how to solve for x

OpenStudy (midhun.madhu1987):

okay.. i'll help

OpenStudy (midhun.madhu1987):

what is x + 2x?

OpenStudy (anonymous):

3x?

OpenStudy (midhun.madhu1987):

yes... the above can be written as: 102 = 2 ( 3x + 6 )

OpenStudy (midhun.madhu1987):

Now divide both sides by 2..

OpenStudy (anonymous):

divide 102 by 2?

OpenStudy (midhun.madhu1987):

yes.. and the right side of equation also

OpenStudy (midhun.madhu1987):

so how will the equation look like?

OpenStudy (anonymous):

51=9x??

OpenStudy (midhun.madhu1987):

51= ( 3x + 6 ) you can add only like terms.. 3x and 6 are not like terms...

OpenStudy (midhun.madhu1987):

Now subtract 6 from both sides.. what u get?

OpenStudy (anonymous):

51-6=45

OpenStudy (midhun.madhu1987):

and what about the right side?

OpenStudy (anonymous):

6-6=0

OpenStudy (midhun.madhu1987):

51 - 6 = 3x + 6 - 6 45 = 3x understood?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is the answer?

OpenStudy (midhun.madhu1987):

now divide both sides by 3 and u get x.

OpenStudy (anonymous):

15

OpenStudy (anonymous):

or 16?